IF A 95% CI FOR THE DIFFERENCE IN TWO INDEPENDENT MEANS IS (-4.5 TO 2.1) THEN THE POINT ESTIMATE IS -2.1.

Answer by stanbon(75887)

IF A 95% CI FOR THE DIFFERENCE IN TWO INDEPENDENT MEANS IS (-4.5 TO 2.1) THEN THE POINT ESTIMATE IS -2.1.
--------------------------------
x-bar = (2.1--4.5)/2 = 6.6/2 = 3.3
======================================
Cheers,
Stan H.